OpenStudy (anonymous):

A triangle has a base of 9 inches and a height of 6 inches. What is the area of the triangle? 36 square inches 56 square inches 54 square inches 27 square inches

OpenStudy (gabebae):

\[A = \frac{ bh }{ 2 }\] that's the area for a triangle

OpenStudy (gabebae):

\[A = \frac{ 9(6) }{ 2 }\] \[A = \frac{ 54 }{ 2 }\] \[A = 27\]